Grade 12Physical Chemistry

If the conc. Of solution of cu/cu^+2 electrod is decreased 10 times then what will be the change in reduction potential??

Milan kansara
8 Years agoGrade 12

1 Answer

Shailendra Kumar Sharma
8 Years ago
Cu+2 +2e → Cu

Eold=Eo + (0.0591/2) log (Cu+2)
with the decrease in concentration by Cu+2/10
Enew=Eo + (0.0591/2) log (Cu+2/10)
       = Eo + (0.0591/2) log (Cu+2)-(0.0591/2)log10
       =Eold – 0.0295